OpenStudy (anonymous):

Anybody here taking Algebra 2 for virtual school and has done lesson the Module 4 Quiz? Can anyone help me with this problem? 4.Mrs. Collins is at the table with you and states that the fourth-degree graphs she has seen have four real zeros. She asks you if it is possible to create a fourth-degree polynomial with only two real zeros. Demonstrate how to do this and explain your steps.

OpenStudy (mathmale):

You could do this by inventing zeros (2 real and 2 complex), writing the corresponding factors, and then multiplying out the four factors. To help you get started: Imagine that your 2 complex zeros are {2-i,2+i,3, 4}. The corresponding factors are (x-2+i),(x-2-i),(x-3), and (x-4). Try multiplying these out to obtain a fourth order polynomial with only two real roots.

OpenStudy (anonymous):

Like multiplying those four together?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Exactly. It's easier to multiply (x-3)(x-4) than to multiply (x-2+i)(x-2-I) (although you still do have to do all of this multiplication). Try multiplying (x-3)(x-4) now. Use FOIL.

OpenStudy (anonymous):

I got x^2-7x+12. @mathmale

OpenStudy (anonymous):

Is that correct? @mathmale

OpenStudy (mathmale):

Great. Now let's begin multiplying out the complex factors. I'd arbitrarily subdivide the info within parentheses as follows: (x-[2+i])(x-[2-i]).

OpenStudy (mathmale):

Hint: use this principle: (a-b)(a+b)=a^2 - b^2. This is a "special product."

OpenStudy (mathmale):

Thus, (x-[2+i])(x-[2-i]) would be written as x^2 - [2+i][2-i]. Have you done this kind of problem before?

OpenStudy (anonymous):

I do not believe so.

OpenStudy (anonymous):

Well at least not including the brackets in a problem.

OpenStudy (mathmale):

Has your teacher (or have your instructional materials) ever discussed "complex numbers"?

OpenStudy (anonymous):

Oh yes she has!

OpenStudy (mathmale):

Good. Let's start with a simpler example: (a+i)(a-i) = a^2 + (what?)

OpenStudy (anonymous):

Would I foil in this case?

OpenStudy (mathmale):

Recall that (a-b)(a+b)=a^-b^2 when both a and b are real. But here we have (a+i)(a-i), where a is real and i is the imaginary operator. Yes, you would use FOIL to multiply those two binomials together.

OpenStudy (anonymous):

I believe it is a^2+ai+ai-1 I am pretty sure I am incorrect but I am not sure.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Hello?@mathmale